# StructEval: Benchmarking LLMs’ Capabilities to Generate Structural Outputs

### v1

@@ 5.3 Structured Output Generation
The ability to generate structured outputs is central to many real-world applications of LLMs (Gu et al., 2024; Tang et al., 2023). These outputs are not only expected to be semantically coherent but must also adhere strictly to syntactic and structural constraints—violations of which can lead to parsing failures, rendering errors, or broken downstream applications. Common tasks include generating JSON for API responses (Geng et al., 2025), YAML or TOML for configuration files (Peddireddy, 2024), HTML or React for UI components (Si et al., 2024), and LaTeX or Markdown for technical writing (Wen et al., 2024). Moreover, in data science, models are used to transform unstructured descriptions into structured formats like CSV or tables for integration into analysis pipelines (Li et al., 2023; Su et al., 2024). In publishing and education, tools that convert textual prompts into diagrams (e.g., using TikZ, SVG, or Mermaid) help automate visualization generation (Lee et al., 2025; Rodriguez et al., 2025; Ku et al., 2025). Despite its significance, structured output generation remains challenging due to the need for models to internalize both syntax rules and hierarchical schema relationships across a wide variety of formats. Our StructEval first conducts a comprehensive evaluation of existing LLMs on both renderable and non-renderable tasks, showing that they still struggle to correctly generate some data formats including TOML, SVG, and Mermaid.
@@ A.4 Task Generation Prompt
Sample Prompt You are a prompt-design assistant building benchmark items for conversion tasks. Input Format: {input_type} Output Format: {output_type} Your task: Think silently through the checklist and then output a single JSON object with: "raw_output_metric": dot-paths for the expected keys/attributes in the {output_type} structure "query": A generated input format {input_type} code inside <code>...</code> tags. Assumed Mapping Rule (state it implicitly in the paths): No XML attributes unless absolutely necessary.

### v2

@@ 5.3 Structured Output Generation
The ability to generate structured outputs is central to many real-world applications of LLMs (gu2024structext; tang-etal-2024-struc). These outputs are not only expected to be semantically coherent but must also adhere strictly to syntactic and structural constraints—violations of which can lead to parsing failures, rendering errors, or broken downstream applications. Common tasks include generating JSON for API responses (Geng2025JSONSchemaBenchAR), YAML or TOML for configuration files (Peddireddy2024EffectiveWA), HTML or React for UI components (Si2024Design2CodeBM), and LaTeX or Markdown for technical writing (Wen2024OverleafCopilotEA). Moreover, in data science, models are used to transform unstructured descriptions into structured formats like CSV or tables for integration into analysis pipelines (Li2023TableGPTTG; Su2024TableGPT2AL). In publishing and education, tools that convert textual prompts into diagrams (e.g., using TikZ, SVG, or Mermaid) help automate visualization generation (Lee2025FromTT; Rodriguez2023StarVectorGS; Ku2025TheoremExplainAgentTM). Despite its significance, structured output generation remains challenging due to the need for models to internalize both syntax rules and hierarchical schema relationships across a wide variety of formats. Our StructEval first conducts a comprehensive evaluation of existing LLMs on both renderable and non-renderable tasks, showing that they still struggle to correctly generate some data formats including TOML, SVG, and Mermaid.
@@ 6 Conclusion
Acknowledgement
This research was supported in part by the Google Cloud Research Credits Program.
@@ A.4 Task Generation Prompt
Sample Prompt You are a prompt-design assistant building benchmark items for conversion tasks. Input Format: {input_type}
Output Format: {output_type}
Your task: Think silently through the checklist and then output a single JSON object with: "raw_output_metric": dot-paths for the expected keys/attributes in the {output_type} structure "query": A generated input format {input_type} code inside <code>...</code> tags. Assumed Mapping Rule (state it implicitly in the paths): No XML attributes unless absolutely necessary.